OpenStudy (anonymous):

Help much appreciated!!!!!! Just a simple algebra 2 question!

OpenStudy (anonymous):

The arc length of a circle is equal to the radius of the circle times the angle, \[L=r\theta=(5 cm)(\pi/3)\]

OpenStudy (anonymous):

Yep, it's D :)

OpenStudy (anonymous):

Glad I could help!

OpenStudy (anonymous):

Sure, I'll do my best

OpenStudy (anonymous):

For g(x), the minimum is -2, because this is the y value where the graph reaches a minimum. For h(x), the minimum is 3, since this is the smallest y value in the table. Since these two minima aren't the same, it can't be A. What do you think about the minimum of f(x)?

OpenStudy (anonymous):

For 12, you want the maximum and minimum values. Remember that the sine function can go from -1 to 1. So to get the maximum value, let the sine function be equal to 1. The result is going to be 35, which is the maximum. To get the minimum, let the sine function be -1. Now you get -15 which is the minimum. To get the period, you take 2pi and divide it by the term inside the sine function that is next to the x. So the period is (2pi)/(pi/2) which is 4. So your answer would be A

OpenStudy (anonymous):

As for 13, you simply plug in t=4 to the equation and evaluate it

OpenStudy (anonymous):

I'm getting C, 4.5: \[5\cos(\pi/6 * 4)+7=5\cos(2\pi/3)+7=4.5\]

OpenStudy (anonymous):

I agree :)

OpenStudy (anonymous):

@vinnv226

OpenStudy (anonymous):

Unfortunately no, you'd need to ask each question independently. But glad I could help

OpenStudy (anonymous):

Thanks for the heads up!!